10.-/1 pointsTGinterAlgHS 5.7.048.CMI Factor the polynomial. m2 n2+4m +4n Read I Watch It Talk to a Tutor 11.-/1 pointsTGInterAlgH5 5.7.055.TP Factor the sum of cubes. 343m3 m-n)(m O (7m-n) (7m2-Trnn + n2) o (7m -n) (49m2 -7mn +n2) mn - Need Help? Redit Talk to a Tuter 12.-/1 pointsTGInterAlgH5 5.7.059. Factor the difference of cubes. x3-Bys
Answer #1

10) Given m^2-n^2+4m+4n

= (m-n)(m+n)+4(m+n)

= [(m-n)+4](m+n)

Therefore, the factors of m^2-n^2+4m+4n are (m-n+4) and (m+n).

11) Given 343m^3+n^3

= (7m)^3+n^3

= (7m+n)[(7m)^2-7m*n+n^2]

= (7m+n)(49m^2-7mn+n^2)

12) Given x^3-8y^6

= x^3-(2y^2)^3

= (x-2y^2)[x^2+x*2y^2+(2y^2)^2]

= (x-2y^2)(x^2+2xy^2+4y^4)

The factors of x^3-8y^6 are (x-2y^2) and (x^2+2xy^2+4y^4) .